# Modeling $(\infty,1)$-categories with Segal spaces

### v1

@@ 3.6. Trivial fibrations are weak equivalences
Since $j_{X}$ and $j_{Y}$ are $\mathcal{J}_{0}$-cofibrations, the induced maps $(j_{X})_{0}\colon X_{0}\to\widetilde{X}_{0}$ and $(j_{Y})_{0}\colon Y_{0}\to\widetilde{Y}_{0}$ are weak equivalences in $\mathrm{sSet}$ by Lemma 3.13. We therefore obtain levelwise weak equivalences $\mathrm{cosk}_{0}((j_{X})_{0})$ and $\mathrm{cosk}_{0}((j_{Y})_{0})$. In the right diagram above, the right-hand square is therefore levelwise a homotopy pullback in $\mathrm{sSet}$. The left-hand square is also levelwise a homotopy pullback in $\mathrm{sSet}$: the right vertical map is levelwise a fibration in $\mathrm{sSet}$ (which is right proper) and $X\to Y\times_{\mathrm{cosk}_{0}(Y_{0})}\mathrm{cosk}_{0}(X_{0})$ is a levelwise weak equivalence in $\mathrm{sSet}$ by Lemma 2.11 (iii). It follows that the composite of the two squares is levelwise a homotopy pullback in $\mathrm{sSet}$, i.e., the map
@@
where $P$ and $Q$ are defined by taking pullbacks. Because $\widetilde{f}\colon\widetilde{X}\to\widetilde{Y}$ is a $\mathcal{J}_{0}$-fibration between $\mathcal{J}_{0}$-fibrant objects, Remarks 3.10 and 3.13 show that the map $\widetilde{f}_{0}\colon\widetilde{X}_{0}\to\widetilde{Y}_{0}$ is a (Kan) fibrant replacement of the fibration $f_{0}$ in $\mathrm{sSet}$. Since $f_{0,0}$ is surjective, so are the isomorphic maps $\pi_{0}(f_{0})$ and $\pi_{0}(\widetilde{f}_{0})$. Since $\widetilde{f}_{0}$ a fibration in $\mathrm{sSet}$, the map $\widetilde{f}_{0,0}$ is then surjective as well. Lemma 3.19 therefore implies that $\mathrm{cosk}_{0}(\widetilde{f}_{0})$ is an $\mathcal{I}$-fibration, so that its pullback $Q\to\widetilde{Y}$ is an $\mathcal{I}$-fibration between Segal spaces. Consequently, it is a Dwyer–Kan equivalence by Proposition 3.17.
In addition, Remarks 3.10 and 3.18 imply that $P\to Q$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$, as it is the pullback of the $\mathcal{J}_{0}$-cofibration $j_{Y}\colon Y\to\widetilde{Y}$ along $\mathrm{cosk}_{0}(\widetilde{f}_{0})$. The map $X\to P$ is a levelwise weak equivalence by the above argument and the map $j_{X}\colon X\to\widetilde{X}$ is a $\mathcal{J}_{0}$-cofibration, and so a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ by Remark 3.10. It follows from $2$-out-of-$3$ that $\widetilde{X}\to Q$ is a weak equivalence between Segal spaces in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$. By Example 2.12, the map $\widetilde{X}\to Q$ is then a Dwyer–Kan equivalence, so that the composite $\widetilde{f}\colon\widetilde{X}\to\widetilde{Y}$ is a Dwyer–Kan equivalence. ∎
@@ 4. Quillen equivalences with models of $(\infty,1)$-categories
In this section, we show that the categorical model structure $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ constructed in Theorem 3.1 is a model of $(\infty,1)$-categories. Specifically, in Section 4.1, we prove that it is Quillen equivalent to the complete Segal space model structure on $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ via the identity adjunction. Additionally, in Section 4.2, we compare it with the model of $(\infty,1)$-categories given by the Segal categories. Denoting by $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ the full subcategory of $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ consisting of the simplicial spaces $X$ such that $X_{0}$ is a set, we show that the inclusion $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\hookrightarrow\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is both a left and a right Quillen equivalence when $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ is endowed with the Segal category model structure. As a consequence, we get that the Segal category model structure is both left- and right-induced from the categorical model structure $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$.
@@ 4.1. Quillen equivalence with complete Segal spaces
We begin by showing that the categorical model structure that we constructed in Theorem 3.1 is Quillen equivalent to Rezk’s model for $(\infty,1)$-categories, given by the complete Segal spaces. To this end, we first review their definition, as well as their associated model structure.
@@
    the fibrant objects are the complete Segal space,
@@
Let $j_{Y}\colon Y\to\widetilde{Y}$ be a $\mathcal{J}$-fibrant replacement and factor the composite $j_{Y}f$ into a $\mathcal{J}$-cofibration $j_{X}\colon X\to\widetilde{X}$ followed by a $\mathcal{J}$-fibration $\tilde{f}\colon\widetilde{X}\to\widetilde{Y}$. Since $\mathcal{J}$-cofibrations are trivial cofibrations in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$, it follows by $2$-out-of-$3$ that $\tilde{f}$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ between Segal spaces. By Theorem 4.2 (iii), the map $\tilde{f}$ is a Dwyer–Kan equivalence between Segal spaces, and hence a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. As $\mathcal{J}$-cofibrations are in particular weak equivalences in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, the map $f$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ by the $2$-out-of-$3$ property. ∎
@@
[…]
@@
By Proposition 3.8, the functor $\mathrm{id}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ preserves cofibrations. By Proposition 3.12, the functor $\mathrm{id}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ preserves fibrations between fibrant objects. Hence the identity adjunction is a Quillen pair by [Joy08, Proposition E.2.14].
Next, we show that the derived unit is a weak equivalence. Let $X$ be a cofibrant object in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ and consider a fibrant replacement $X\to\widehat{X}$ in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. Then by Lemma 4.3 the component of the derived unit $X\to\widehat{X}$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$.
Finally, we show that the derived counit is a weak equivalence. Let $X$ be a fibrant object in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ and consider a cofibrant replacement $\overline{X}\to X$ in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. We want to prove that the component of the derived counit $\overline{X}\to X$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. Without loss of generality, we can assume that $\overline{X}\to X$ is a trivial fibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, i.e., an $\mathcal{I}$-fibration. Since the complete Segal space $X$ is in particular a Segal space, then so is $\overline{X}$. Hence, by Proposition 3.17, the map $\overline{X}\to X$ is a Dwyer–Kan equivalence between Segal spaces, and so it is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ by Theorem 4.2 (iii). ∎
@@ 4.2. Quillen equivalences with Segal categories
    the fibrant objects are the Segal spaces,
@@
The Segal category model structure is also a model of $(\infty,1)$-categories, as it is Quillen equivalent to the complete Segal space model structure by [Ber07b, Theorem 6.3]. This Quillen equivalence is induced by the inclusion $I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\hookrightarrow\mathrm{sSet}^{\Delta^{\mathrm{op}}}$, which we recall admits both a left adjoint $L\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}\to\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ and a right adjoint $R\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}\to\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$, see Remark 2.9.
@@
[…]
@@
In the categorical model structure that we constructed in Theorem 3.1, we removed all the problematic monomorphisms whose image under $L$ is not a monomorphism. As a consequence, the adjunction $L\dashv I$ does become a Quillen equivalence between $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ and $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. We prove the following.
@@
[…]
@@
[…]
@@
By [Joy08, Proposition E.2.14], it remains to show that $L\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\to\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ sends maps in $\mathcal{J}$ to weak equivalences in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. First, consider a map $X\hookrightarrow Y$ in $\mathcal{J}_{0}$, i.e., a map of the form (i) or (ii) in Notation 3.9. Using Remark 4.6, we see that the fibrant replacement of the induced map $LX\hookrightarrow LY$ can be computed to be the identity, and so the map $LX\hookrightarrow LY$ is a weak equivalence in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. Finally, if $X\hookrightarrow Y$ is the map $F[0]\hookrightarrow NI[1]$, then it is a Dwyer–Kan equivalence between Segal spaces and so it is also a weak equivalence in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. ∎
@@
[…]
@@
Since $I$ is a right adjoint, it preserves monomorphisms. Hence, it remains to show that, for every monomorphism $X\hookrightarrow Y$ in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$, the monomorphism $IX\hookrightarrow IY$ satisfies condition (2) of Proposition 3.8. But this follows directly from the fact that $X_{0}$ and $Y_{0}$ are sets. This shows that $I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ preserves cofibrations.
We now prove that $I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ preserves weak equivalences. By Remark 4.6, we see that $I$ preserves fibrant replacements. Hence, it suffices to show that $I$ preserves weak equivalences between fibrant objects, but this is clear as in both model structures these are given by the Dwyer–Kan equivalences between Segal spaces. ∎
To prove that the above are further Quillen equivalences, we show that the following diagram of right Quillen functors induces a commutative diagram at the level of homotopy categories. We then deduce from Theorems 4.4 and 4.7 and the $2$-out-of-$3$ property for Quillen equivalences that $I$ is a Quillen equivalence, thereby concluding the proof of Theorem 4.9.
@@
Let $X$ be a fibrant object in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$, i.e., a complete Segal space. Since all objects in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ are cofibrant, the component of the counit $\varepsilon_{X}\colon IRX\to X$ coincides with the component of the derived counit at $X$. As $IRX$ and $X$ are both Segal spaces and $\varepsilon_{X}$ induces a bijection $(IRX)_{0,0}\cong X_{0,0}$, it follows directly from Remarks 2.9 and 2.11 that $\varepsilon_{X}\colon IRX\to X$ is a Dwyer–Kan equivalence, and so a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. ∎
@@
We recover Corollary 4.13 as an application of the following abstract argument, by taking $\mathcal{M}=\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$, $\mathcal{N}=\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, and $F=I$.
@@
Since $F\colon\mathcal{M}\to\mathcal{N}$ is both left and right Quillen, it preserves (trivial) fibrations and (trivial) cofibrations. Every weak equivalence factors as a trivial cofibration followed by a trivial fibration, so that $F$ preserves all weak equivalences. It remains to show that $F$ also reflects cofibrations, fibrations, and weak equivalences. We will only treat the fibrations and weak equivalences, the case of cofibrations following dually.
Since $F\colon\mathcal{M}\to\mathcal{N}$ preserves fibrant replacements and the derived counit of the adjunction formed by $F$ and its left adjoint is a weak equivalence in $\mathcal{M}$, it follows that $F$ reflects all weak equivalences. To see that $F$ reflects fibrations, let $X\to Y$ be a map in $\mathcal{M}$ such that the induced map $FX\twoheadrightarrow FY$ is a fibration in $\mathcal{N}$. To see that $X\to Y$ is a fibration in $\mathcal{M}$, we need to show that it has the right lifting property against every trivial cofibration $A\xhookrightarrow{\simeq}B$ in $\mathcal{M}$. Since $F$ is fully faithful, this is equivalent to showing that $FX\twoheadrightarrow FY$ has the right lifting property against $FA\to FB$, for every trivial cofibration $A\xhookrightarrow{\simeq}B$ in $\mathcal{M}$. This follows from the fact that $F$ preserves trivial cofibrations. ∎

### v2

@@ 3.6. Trivial fibrations are weak equivalences
Since $j_{X}$ and $j_{Y}$ are $\mathcal{J}_{0}$-cofibrations, the induced maps $(j_{X})_{0}\colon X_{0}\to\widetilde{X}_{0}$ and $(j_{Y})_{0}\colon Y_{0}\to\widetilde{Y}_{0}$ are weak equivalences in $\mathrm{sSet}$ by Lemma˜3.13. We therefore obtain levelwise weak equivalences $\mathrm{cosk}_{0}((j_{X})_{0})$ and $\mathrm{cosk}_{0}((j_{Y})_{0})$. In the right diagram above, the right-hand square is therefore levelwise a homotopy pullback in $\mathrm{sSet}$. The left-hand square is also levelwise a homotopy pullback in $\mathrm{sSet}$: the right vertical map is levelwise a fibration in $\mathrm{sSet}$ (which is right proper) and $X\to Y\times_{\mathrm{cosk}_{0}(Y_{0})}\mathrm{cosk}_{0}(X_{0})$ is a levelwise weak equivalence in $\mathrm{sSet}$ by Lemma 3.19 (ii). It follows that the composite of the two squares is levelwise a homotopy pullback in $\mathrm{sSet}$, i.e., the map
@@
where $P$ and $Q$ are defined by taking pullbacks. Because $\widetilde{f}\colon\widetilde{X}\to\widetilde{Y}$ is a $\mathcal{J}_{0}$-fibration between $\mathcal{J}_{0}$-fibrant objects, Remarks˜3.10 and 3.13 show that the map $\widetilde{f}_{0}\colon\widetilde{X}_{0}\to\widetilde{Y}_{0}$ is a (Kan) fibrant replacement of the fibration $f_{0}$ in $\mathrm{sSet}$. Since $f_{0,0}$ is surjective, so are the isomorphic maps $\pi_{0}(f_{0})$ and $\pi_{0}(\widetilde{f}_{0})$. Since $\widetilde{f}_{0}$ a fibration in $\mathrm{sSet}$, the map $\widetilde{f}_{0,0}$ is then surjective as well. Lemma˜3.19 (i) therefore implies that $\mathrm{cosk}_{0}(\widetilde{f}_{0})$ is an $\mathcal{I}$-fibration, so that its pullback $Q\to\widetilde{Y}$ is an $\mathcal{I}$-fibration between Segal spaces. Consequently, it is a Dwyer–Kan equivalence by Proposition˜3.17.
In addition, Remarks˜3.10 and 3.18 imply that $P\to Q$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$, as it is the pullback of the $\mathcal{J}_{0}$-cofibration $j_{Y}\colon Y\to\widetilde{Y}$ along $\mathrm{cosk}_{0}(\widetilde{f}_{0})$. The map $X\to P$ is a levelwise weak equivalence by the above argument and the map $j_{X}\colon X\to\widetilde{X}$ is a $\mathcal{J}_{0}$-cofibration, and so a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ by Remark˜3.10. It follows from 22-out-of-33 that $\widetilde{X}\to Q$ is a weak equivalence between Segal spaces in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$. By Example˜2.12, the map $\widetilde{X}\to Q$ is then a Dwyer–Kan equivalence, so that the composite $\widetilde{f}\colon\widetilde{X}\to\widetilde{Y}$ is a Dwyer–Kan equivalence. ∎
@@ 4. Quillen equivalences with models of $(\infty,1)$-categories
In this section, we show that the categorical model structure $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ constructed in Theorem˜3.1 is a model of $(\infty,1)$-categories. Specifically, in Section˜4.1, we prove that it is Quillen equivalent to the complete Segal space model structure on $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ via the identity adjunction. Additionally, in Section˜4.2, we compare it with the model of $(\infty,1)$-categories given by the Segal categories. Denoting by $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ the full subcategory of $\mathrm{sSet}^{\Delta^{\mathrm{op}}}$ consisting of the simplicial spaces $X$ such that $X_{0}$ is a set, we show that the inclusion $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\hookrightarrow\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ is both a left and a right Quillen equivalence when $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ is endowed with the Segal category model structure. As a consequence, we get that the Segal category model structure is both left- and right-induced from the categorical model structure $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$.
@@ 4.1. Quillen equivalence with complete Segal spaces
We begin by showing that the categorical model structure that we constructed in Theorem˜3.1 is Quillen equivalent to Rezk’s model for $(\infty,1)$-categories, given by the complete Segal spaces. To this end, we first review their definition, as well as their associated model structure.
@@
    the fibrant objects are the complete Segal spaces,
@@
Let $j_{Y}\colon Y\to\widetilde{Y}$ be a $\mathcal{J}$-fibrant replacement and factor the composite $j_{Y}f$ into a $\mathcal{J}$-cofibration $j_{X}\colon X\to\widetilde{X}$ followed by a $\mathcal{J}$-fibration $\tilde{f}\colon\widetilde{X}\to\widetilde{Y}$. Since $\mathcal{J}$-cofibrations are trivial cofibrations in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$, it follows by 22-out-of-33 that $\tilde{f}$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ between Segal spaces. By Theorem˜4.2 (iii), the map $\tilde{f}$ is a Dwyer–Kan equivalence between Segal spaces, and hence a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. As $\mathcal{J}$-cofibrations are in particular weak equivalences in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, the map $f$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ by the 22-out-of-33 property. ∎
@@
$\mathrm{id}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\rightleftarrows\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0mu\mathrm{id}$
@@
By Proposition˜3.8, the functor $\mathrm{id}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ preserves cofibrations. By Proposition˜3.12, the functor $\mathrm{id}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ preserves fibrations between fibrant objects. Hence the identity adjunction is a Quillen pair by [Joy08, Proposition E.2.14].
Next, we show that the derived unit is a weak equivalence. Let $X$ be a cofibrant object in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ and consider a fibrant replacement $X\to\widehat{X}$ in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. Then by Lemma˜4.3 the component of the derived unit $X\to\widehat{X}$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$.
Finally, we show that the derived counit is a weak equivalence. Let $X$ be a fibrant object in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ and consider a cofibrant replacement $\overline{X}\to X$ in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ given by an $\mathcal{I}$-fibration. We want to prove that the component of the derived counit $\overline{X}\to X$ is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$. Since the complete Segal space $X$ is in particular a Segal space, then so is $\overline{X}$. Hence, by Proposition˜3.17, the map $\overline{X}\to X$ is a Dwyer–Kan equivalence between Segal spaces, and so it is a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$ by Theorem˜4.2 (iii). ∎
@@ 4.2. Quillen equivalences with Segal categories
    the fibrant objects are the Segal categories, i.e., Segal spaces with a discrete space of objects,
@@
The Segal category model structure is also a model of $(\infty,1)$-categories, as it is Quillen equivalent to the complete Segal space model structure by [Ber07b, Theorem 6.3]. This Quillen equivalence is induced by the inclusion $I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\hookrightarrow\mathrm{sSet}^{\Delta^{\mathrm{op}}}$, which we recall admits both a left adjoint $L\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}\to\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ and a right adjoint $R\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}\to\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$, see Remark˜2.9.
@@
The adjunction $I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\rightleftarrows\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0muR$ is a Quillen equivalence.
@@
In the categorical model structure that we constructed in Theorem˜3.1, we removed all the problematic monomorphisms whose image under $L$ is not a monomorphism. As a consequence, the adjunction $L\dashv I$ does become a Quillen equivalence between $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ and $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. We prove the following.
@@
$L\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\rightleftarrows\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0muI\quad\text{and}\quad I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\rightleftarrows\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0muR$
@@
The adjunction $L\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\rightleftarrows\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0muI$ is a Quillen pair.
@@
By [Joy08, Proposition E.2.14], it remains to show that $L\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\to\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ sends maps in $\mathcal{J}$ to weak equivalences in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. First, consider a map $X\hookrightarrow Y$ in $\mathcal{J}_{0}$, i.e., a map of the form (i) or (ii) in Notation˜3.9. Using Remark˜4.6, we see that the fibrant replacement of the induced map $LX\hookrightarrow LY$ can be computed to be the identity, and so the map $LX\hookrightarrow LY$ is a weak equivalence in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. Finally, if $X\hookrightarrow Y$ is the map $F[0]\hookrightarrow NI[1]$, then it is a Dwyer–Kan equivalence between Segal spaces and so it is also a weak equivalence in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. ∎
@@
The adjunction $I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\rightleftarrows\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0muR$ is a Quillen pair.
@@
Suppose that $X\hookrightarrow Y$ is a monomorphism in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$. Since $I$ is a right adjoint and $X_{0}$ and $Y_{0}$ are sets, the monomorphism $IX\hookrightarrow IY$ satisfies condition (2) of Proposition˜3.8, hence it is a cofibration. This shows that $I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ preserves cofibrations.
We now prove that $I\colon\mathcal{P}\mathrm{Cat}(\mathrm{sSet})\to\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$ preserves weak equivalences. By Remark˜4.6, we see that $I$ preserves fibrant replacements. Hence, it suffices to show that $I$ preserves weak equivalences between fibrant objects, but this is clear as in both model structures these are given by the Dwyer–Kan equivalences between Segal spaces. ∎
To prove that the above are further Quillen equivalences, we show that the following diagram of right Quillen functors induces a commutative diagram at the level of homotopy categories. We then deduce from Theorems˜4.4 and 4.7 and the 22-out-of-33 property for Quillen equivalences that $I$ is a Quillen equivalence, thereby concluding the proof of Theorem˜4.9.
@@
Let $X$ be a fibrant object in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{CSS}}$, i.e., a complete Segal space. Since all objects in $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ are cofibrant, the component of the counit $\varepsilon_{X}\colon IRX\to X$ coincides with the component of the derived counit at $X$. As $IRX$ and $X$ are both Segal spaces and $\varepsilon_{X}$ induces a bijection $(IRX)_{0,0}\cong X_{0,0}$, it follows directly from Remarks˜2.9 and 2.11 that $\varepsilon_{X}\colon IRX\to X$ is a Dwyer–Kan equivalence, and so a weak equivalence in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$. ∎
@@
We recover Corollary˜4.13 as an application of the following abstract argument, by taking $\mathcal{M}=\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$, $\mathcal{N}=\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Cat}}$, and $F=I$.
@@
Since $F\colon\mathcal{M}\to\mathcal{N}$ is both left and right Quillen, it preserves (trivial) fibrations and (trivial) cofibrations. We further prove that it reflects (trivial) fibrations and (trivial) cofibrations. Then, since a map is a weak equivalence if and only if it factors as a trivial cofibration followed by a trivial fibration, then $F$ also preserves and reflects all weak equivalences.
We only show that $F$ reflects fibrations; the case of trivial fibrations is similar, and the case of (trivial) cofibrations is dual. Let $X\to Y$ be a map in $\mathcal{M}$ such that the induced map $FX\twoheadrightarrow FY$ is a fibration in $\mathcal{N}$. To see that $X\to Y$ is a fibration in $\mathcal{M}$, we need to show that it has the right lifting property against every trivial cofibration $A\xhookrightarrow{\simeq}B$ in $\mathcal{M}$. Since $F$ is fully faithful, this is equivalent to showing that $FX\twoheadrightarrow FY$ has the right lifting property against $FA\to FB$, for every trivial cofibration $A\xhookrightarrow{\simeq}B$ in $\mathcal{M}$. This follows from the fact that $F$ preserves trivial cofibrations. ∎